OpenStudy (friendzone):

Math Help A two-sided coin (heads/tails) is flipped once, and one card is drawn from a standard 52 deck of playing cards. Calculate the probability of each given situation, and drag the correct answer to it. Flipping a Heads AND drawing a Heart Flipping a Heads OR drawing a Heart Flipping a Heads Drawing a Heart NOT drawing a King Flipping a Tails AND drawing a King Flipping a Tails OR drawing a King Drawing a King OR drawing a Heart Flipping a Heads OR flipping a Tails Drawing a King AND drawing a Heart 1 1/4 1/52 1/8 12/13 15/26 3/4 4/13 1/2 1/26

OpenStudy (anonymous):

Would you like to know how to do this or are only concerned with the answer?

OpenStudy (friendzone):

@1DEA i know how just checking my work like always ..

OpenStudy (anonymous):

ahh ok, I wasn't sure because if I had to throughly explain to how to get all the answers...it would take a long time. Do you want me just to explain each one, or do you have a particular one you don't understand that well?

OpenStudy (friendzone):

no i just hope my Answers are right lol just put = too the ones lol

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

Flipping a Heads AND drawing a Heart = 1/8 Flipping a Heads OR drawing a Heart = 3/4 Flipping a Heads = 1/2 Drawing a Heart =1/4 NOT drawing a King = 12/13 Flipping a Tails AND drawing a King = 1/26 Flipping a Tails OR drawing a King = 15/26 Drawing a King OR drawing a Heart = 4/13 Flipping a Heads OR flipping a Tails = 1 Drawing a King AND drawing a Heart = 1/52 sorry for taking so long, mental math problems lol. I think that's correct let me know if you have any disagreements.

OpenStudy (friendzone):

flipping tails and drawing a king =1/26 ?

OpenStudy (friendzone):

i think its 1/52 switch those two i think

OpenStudy (anonymous):

the probability of flipping a tail is 1/2, right? and the probability of drawing a king is 1/13 since you want both you multiply the two making 1/26.

OpenStudy (friendzone):

oh snap that fooled me i hate this kind of stuff

OpenStudy (anonymous):

yup probability is by far the hardest Math, even more than linear, discrete, or multivariate

OpenStudy (friendzone):

thanks lol i got some more stuff too lol

OpenStudy (anonymous):

if you want me to help sure it's not a problem.

OpenStudy (friendzone):

This is the ratio of the number of ways the event can occur to the number of ways the event cannot occur. This is the number of selected outcomes divided by the total number of possible outcomes. It is a number between 0 and 1, including 0 and 1. This is an outline designed to demonstrate or explain the number of times a specified periodic phenomenon occurs within a specified interval. The desired outcomes of a specified event. The different possible results from a probability model. This helps to visually display the outcomes of an experiment consisting of a series of activities (rolling dice multiple times, total pizza choices, etc.). The total number of outcomes corresponds to the total number of final branches in the diagram. An outcome in a probability experiment. Events in which the outcome of one event affects the outcome of the other event. A diagram that uses circles or ovals to illustrate the relationship between sets. The mathematical calculation that an event will happen. The ratio of the number of times an event occurs to the total number of trials. This is the probability that event B will occur given that event A has occurred. The sum of the probabilities of each outcome multiplied by the outcome value. For probability this rule states that (A ∪ B) = P(A) + P(B) - P(A ∩ B) A complementary event is the event that doesn't happen. If the probability of an event occurring is A, the the probability of the complementary event is 1-A. The sum of the probability of an event and the complementary event is Dependent Events Event Favorable Outcomes Frequency Diagram OddsOutcomes Probability Tree Diagram Venn Diagramaddition rulecomplementary eventsconditional probability expected value experimental probability theoretical probability

OpenStudy (friendzone):

Dependent events Event Favorable outcomes Frequency diagram Odds outcomes Probability Tree Diagram Venn Diagram Addition Rule conditional probability complementary events Expected value experimental probability theoretical probability

OpenStudy (friendzone):

had to put it in order its like matching

OpenStudy (anonymous):

that's a huge wall of text lol...let me read it

OpenStudy (anonymous):

Oh I see got it matching...

OpenStudy (friendzone):

its at each period :)

OpenStudy (anonymous):

so you want the same as before?

OpenStudy (friendzone):

yea lol my bad

OpenStudy (anonymous):

This is the ratio of the number of ways the event can occur to the number of ways the event cannot occur. - Odds This is the number of selected outcomes divided by the total number of possible outcomes. It is a number between 0 and 1, including 0 and 1. - Theoretical Probability This is an outline designed to demonstrate or explain the number of times a specified periodic phenomenon occurs within a specified interval. -Frequency diagram The desired outcomes of a specified event. - Favorable outcomes The different possible results from a probability model. -outcomes This helps to visually display the outcomes of an experiment consisting of a series of activities (rolling dice multiple times, total pizza choices, etc.). The total number of outcomes corresponds to the total number of final branches in the diagram. - Tree Diagram An outcome in a probability experiment. - Event Events in which the outcome of one event affects the outcome of the other event. - Dependent events A diagram that uses circles or ovals to illustrate the relationship between sets. - Venn Diagram The mathematical calculation that an event will happen. - Probability The ratio of the number of times an event occurs to the total number of trials. - Experimental Probability This is the probability that event B will occur given that event A has occurred. - conditional probability The sum of the probabilities of each outcome multiplied by the outcome value. - Expected value For probability this rule states that (A B) = P(A) + P(B) - P(A B) - (most likely the addition rule, can't really tell due to the question marks) A complementary event is the event that doesn't happen. If the probability of an event occurring is A, the the probability of the complementary event is 1-A. The sum of the probability of an event and the complementary event is - 1???? The remaining answer is complementary event but honestly it doesn't make much Oh also I am uncertain about whether theoretical probability or probability should be switched

OpenStudy (anonymous):

*make much sense.

OpenStudy (friendzone):

thanks lol @1DEA